Which sample contains the same number of atoms as 24 grams of carbon? (1) 10. g Ne (2) 24g Mg (3) 80. g Ar (4) 4.0 g He (Can you explain or write it out please)


1
Expert's answer
2021-03-09T06:16:45-0500

M(C) = 12 g mol-1;

M(Ne) = 20 g mol-1;

M(Mg) = 24 g mol-1;

M(Ar) = 40 g mol-1;

M(He) = 4 g mol-1.


Solution:

One mole of any substance contains 6.022×1023 atoms/molecules (Na).

Hence,

Number of atoms in 24 g of C = (24 g / 12 g mol-1) × Na = 2.0 × Na atoms


(1): Number of atoms in 10 g of Ne = (10 g / 20 g mol-1) × Na = 0.5 × Na atoms

(2): Number of atoms in 24 g of Mg = (24 g / 24 g mol-1) × Na = 1.0 × Na atoms

(3): Number of atoms in 80 g of Ar = (80 g / 40 g mol-1) × Na = 2.0 × Na atoms

(4): Number of atoms in 4 g of He = (4 g / 4 g mol-1) × Na = 1.0 × Na atoms


Hence, (3) (80 grams of argon) contains same number of atoms (2Na atoms) as in 24 grams of carbon.


Answer: 80 grams of argon (Ar) contains same number of atoms (2.0 × Na atoms) as in 24 grams of carbon (C).